OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (anonymous):

need help n explain it to me

OpenStudy (anonymous):

g(x) = 2(x-4) 20 = 2x-8 20+8=2x 28=2x 28/2=x 14=x

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

help me sir

OpenStudy (lgbasallote):

why do you doubt @HeyItzAri16 's answer?

OpenStudy (anonymous):

because i think it's 32

OpenStudy (lgbasallote):

will you show your solution?

OpenStudy (lgbasallote):

oh wait nevermind

OpenStudy (lgbasallote):

i saw what you did

OpenStudy (anonymous):

ight 2(20-4) 2(16) 32

OpenStudy (lgbasallote):

g(x) = 20 means you equate g(x) which is 2(x-4) to 20... so the equation is 2(x-4) = 20 that is g(x) = 20 what you did...is replaced 20 with the x of g(x)... this is g(20) not g(x) = 20 do you get it?

OpenStudy (anonymous):

yea thxs n soo her answer was right

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

if unsure of your answer, just check it... g(14) = 2(14 - 4) = ....

OpenStudy (anonymous):

if you get what you're s'posed to get then it's correct....